# MRI motion correction via efficient residual-guided denoising diffusion probabilistic models

### v1

@@ 3.2 Quantitative results
Figure 6: Boxplots of PSNR, SSIM, and NMSE metrics across different motion artifact levels.
Table 1: Quantitative Metrics demonstrated in mean±${}_{\pm\text{std}}$ across different motion artifact levels of the in-silico dataset are summarized. The arrows indicate the direction of better performance. Bold indicates the best values.
@@
PNSR [dB] $\uparrow$ Minor $34.62_{\pm 3.25}$ $37.03_{\pm 1.65}$ $41.21_{\pm 2.91}$ $38.25_{\pm 3.08}$ $\bm{41.91_{\pm 2.94}}$
@@
[…]
@@ 3.3 Ablation Study
[…]
Moderate $37.97_{\pm 2.39{\color[rgb]{1,0,0}(2.04)}}$ $0.98_{\pm 0.01{\color[rgb]{1,0,0}(0.48)}}$ $0.24_{\pm 0.16{\color[rgb]{1,0,0}(-14.74)}}$
$\ell_{2}+\ell_{1}$ (Res- MoCoDiff) Heavy $34.15_{\pm 2.42{\color[rgb]{1,0,0}(1.21)}}$ $0.96_{\pm 0.01{\color[rgb]{1,0,0}(0.95)}}$ $0.58_{\pm 0.40{\color[rgb]{1,0,0}(-5.73)}}$
@@ 4 Discussion
In this study, we proposed Res-MoCoDiff, an efficient denoising diffusion probabilistic model designed to reconstruct motion-free images. By leveraging a residual error shifting mechanism (illustrated in Figure 1), our method performs the sampling process in only four steps (see Algorithm 2), thereby facilitating its integration into current clinical practices. Notably, Res-MoCoDiff achieves an average sampling time of 0.37 seconds per batch of two image slices, which is substantially lower than the 101.74 seconds per batch required by the conventional TM-DDPM approach.
@@
Extensive qualitative and quantitative evaluations on both in-silico and in-vivo datasets demonstrate the superior performance of Res-MoCoDiff in removing motion artifacts across different distortion levels. While comparative models often leave residual artifacts—particularly under heavy and moderate motion conditions—Res-MoCoDiff consistently eliminates these imperfections, as highlighted by the green arrows in Figure 4. Moreover, the proposed method excels at recovering fine structural details, resulting in higher pixel-wise correlations (see Figure 4j–l). Although our method achieves the second highest PSNR among the evaluated techniques (see Table 1), its overall improvements in SSIM and NMSE, together with the perceptually superior image quality, underscore its clinical efficacy.
@@
The inclusion of an $\ell_{1}$ regularizer during training further enhances image sharpness by reducing NMSE, particularly important given that higher NMSE is typically associated with blurry reconstructions. This improvement aligns with the increases observed in both PSNR and SSIM, resulting in images that are structurally closer to the ground truth [44].

### v2

@@ 3.2 Quantitative results
Figure 7: Boxplots of PSNR, SSIM, and NMSE metrics across different motion artifact levels for the in-silico IXI dataset.
[…]
@@
PSNR [dB] $\uparrow$ Minor $34.62_{\pm 3.25}$ $37.03_{\pm 1.65}$ $41.21_{\pm 2.91}$ $38.25_{\pm 3.08}$ $\bm{41.91_{\pm 2.94}}$
@@
In the in-vivo MR-ART dataset, the original motion-corrupted images at distortion levels 1 and 2 yield NMSE, SSIM, and PSNR values of $(2.56\pm 2.38\%,$ $0.74{\pm 0.10},$ $28.61{\pm 2.85}\,\text{dB})$ and $(3.30{\pm 2.76}\%,\,0.72\pm 0.10,\,27.51\pm 2.83)$, respectively. Res-MoCoDiff reduces NMSE by $33.21\%$ (to $1.71{\pm 1.49}$) for level 1 and $37.41\%$ (to $2.07{\pm 1.79}$) for level 2. Additionally, it raises SSIM by $23.60\%$ (to $0.92\pm 0.05$) for level 1 and $25.58\%$ (to $0.91\pm 0.05$) for level 2, while also improving PSNR by $6.26\%$ (to $30.40\pm 2.90\,\text{dB}$) for level 1 and $7.71\%$ (to $29.63\pm 2.97\,\text{dB}$) for level 2. These gains underline the robust performance of Res-MoCoDiff for both in-silico and in-vivo motion artifact correction.
@@ 3.3 Ablation Study
Minor $41.91_{\pm 2.94{\color[rgb]{1,0,0}(3.63\%)}}$ $0.99_{\pm 0.00{\color[rgb]{1,0,0}(0.42\%)}}$ $0.10_{\pm 0.09{\color[rgb]{1,0,0}(-26.29\%)}}$
Moderate $37.97_{\pm 2.39{\color[rgb]{1,0,0}(2.04\%)}}$ $0.98_{\pm 0.01{\color[rgb]{1,0,0}(0.48\%)}}$ $0.24_{\pm 0.16{\color[rgb]{1,0,0}(-14.74\%)}}$
$\ell_{2}+\ell_{1}$ (Res-
MoCoDiff)
Heavy $34.15_{\pm 2.42{\color[rgb]{1,0,0}(1.21\%)}}$ $0.96_{\pm 0.01{\color[rgb]{1,0,0}(0.95\%)}}$ $0.58_{\pm 0.40{\color[rgb]{1,0,0}(-5.73\%)}}$
@@ 4 Discussion
In this study, we proposed Res-MoCoDiff, an efficient denoising diffusion probabilistic model designed to reconstruct motion-free images. By leveraging a residual error shifting mechanism (illustrated in Figure 1), our method performs the sampling process in only four steps (see Algorithm 2), thereby facilitating its integration into current clinical practices. Unlike conventional DDPMs that require hundreds of reverse steps, Res-MoCoDiff leverages residual error shifting to bring the corrupted distribution closer to the motion-free image distribution at the terminal diffusion step. This enables accurate image restoration in only four reverse steps. Our validation experiments showed that using more than four steps did not yield significant improvements, confirming the efficiency of this reduced-step sampling. Notably, Res-MoCoDiff achieves an average sampling time of 0.37 seconds per batch of two image slices, which is substantially lower than the 101.74 seconds per batch required by the conventional TM-DDPM approach.
@@
Extensive qualitative and quantitative evaluations on both in-silico and in-vivo datasets demonstrate the superior performance of Res-MoCoDiff in removing motion artifacts across different distortion levels. While comparative models often leave residual artifacts, particularly under heavy and moderate motion conditions, Res-MoCoDiff consistently eliminates these imperfections, as highlighted by the green arrows in Figure 4. Moreover, the proposed method excels at recovering fine structural details, resulting in higher pixel-wise correlations (see Figure 5(a)-(c)). Although our method achieves the second highest PSNR among the evaluated techniques (see Table 1), its overall improvements in SSIM and NMSE, together with the perceptually superior image quality, underscore its clinical efficacy.
@@
The inclusion of an $\ell_{1}$ regularizer during training further enhanced image sharpness by reducing NMSE, which is particularly important given that higher NMSE is often associated with blurry reconstructions. This improvement was consistent with the increases observed in both PSNR and SSIM, resulting in images that were structurally closer to the ground truth [46].
While $\ell_{2}$ loss is widely used in medical image restoration to reduce global pixel-level error, it can lead to oversmoothed reconstructions. By contrast, the $\ell_{1}$ term preserves high-frequency information and improves edge sharpness, thereby producing more realistic structures. The complementary behavior of $\ell_{1}$ and $\ell_{2}$ has been reported in both general image restoration [47] and MRI artifact removal [48]. Our ablation study (Table 2) confirmed these benefits, showing that the combined loss reduced NMSE while enhancing perceptual quality compared with $\ell_{2}$ alone.
A further consideration in designing Res-MoCoDiff is our decision to operate in the image domain rather than in k-space. While k-space methods can, in principle, provide direct access to raw acquisition information, their use in clinical practice is limited by several practical factors. Raw k-space data are not routinely stored in most hospital archives or large-scale public repositories, and reconstruction pipelines differ substantially across scanner vendors and software versions, which complicates reproducibility. In addition, implementing k-space motion correction often requires vendor-specific software environments that are not widely accessible. By contrast, reconstructed magnitude images are universally available and can be directly integrated into existing clinical workflows. Our image-domain design therefore prioritizes generalizability and clinical feasibility, allowing Res-MoCoDiff to serve as an off-the-shelf solution that can be applied retrospectively across a wide range of studies and institutions.
We focused on T1-w brain MRI because this sequence is both highly susceptible to motion artifacts and is central to Radiation Oncology workflows [49]. This emphasis ensures that our evaluation addresses a clinically important sequence with high impact on both diagnostic and therapeutic decision making.
This study has several limitations that warrant discussion. First, in severely degraded in-vivo cases where the motion-corrupted input lacks soft-tissue contrast, Res-MoCoDiff may hallucinate anatomical details that are not consistent with the ground truth. This limitation reflects the fact that when essential structural information is absent from the input, the model cannot reliably recover it. Future work should therefore explore strategies such as incorporating multi-contrast MRI, embedding stronger physics-informed priors, or using uncertainty quantification to identify high-risk reconstructions. Second, our work was restricted to brain MRI, and other sequences such as diffusion-weighted echo-planar imaging (EPI) and high-resolution T2w imaging were not explored. Motion in brain MRI is typically irregular, unpredictable, and of relatively small amplitude, and the residual-guided design of Res-MoCoDiff is well suited for these conditions. However, larger and non-rigid motion events, such as sneezing or swallowing during long 3D acquisitions, may introduce more complex artifacts that are not fully addressed by the current framework. In addition, our motion simulation followed the established approach of Duffy et al. [12], which reproduces realistic ghosting and ringing but does not explicitly account for the increased probability of motion during longer scans. Furthermore, it should also be clarified that this study addresses retrospective motion artifact correction in structural brain MRI, which is distinct from physiological motion correction approaches (e.g., respiratory or cardiac compensation) used in free-breathing acquisitions. To extend Res-MoCoDiff to these broader scenarios, retraining on sequence-specific data may be required, and future studies could explore whether architectural modifications would further improve performance in modeling non-rigid motion. Addressing these extensions will be important for establishing the generalizability of the framework across diverse imaging protocols and clinical applications. Finally, future work should also explore integration with motion modeling or hybrid prospective-retrospective strategies to improve robustness in more complex scenarios.